Calculate the weight of a machine with mass m = 1.2 t when it passes with velocity v = 20 m / s through the lowest point of a bridge bent down. The radius of the bridge is R = 100 m. Compare this weight with the weight of the car at rest. What is the cause of the change in weight of the car compared to the usual weight?


1
Expert's answer
2021-01-27T14:31:07-0500

The weight of the car at rest can be found as follows:


"W_{rest}=F_N=mg=1.2\cdot10^3\ kg\cdot9.8\ \dfrac{m}{s^2}=11760\ N."

The weight of the car at the lowest point of a bridge bent down can be found as follows:


"W=F_N=\dfrac{mv^2}{R}+mg,""W=F_N=\dfrac{1.2\cdot10^3\ kg\cdot(20\ \dfrac{m}{s})^2}{100\ m}+11760\ N=16560\ N."

Let's compare this weight with the weight of the car at rest:


"\dfrac{W}{W_{rest}}=\dfrac{16560\ N}{11760\ N}=1.4."

As we can see from calculations, the weight of the car at the lowest point of a bridge bent down 1.4 times larger than the weight of the car at rest. The centripetal acceleration cause the change in weight of the car compared to the usual weight.